OpenStudy (anonymous):

Let u = <-5, 1>, v = <7, -4>. Find 9u - 6v.

OpenStudy (kc_kennylau):

Just multiply 9 to the components of u, and 6 to the components of v, and subtract the respective components

OpenStudy (kc_kennylau):

(Am I right to suppose it to be vectors)

OpenStudy (anonymous):

wait i can no idea what your asking

OpenStudy (anonymous):

do you think you can work it out for me? I'm taking an exam that i know nothing about.. ha

OpenStudy (kc_kennylau):

Is it about vectors?

OpenStudy (anonymous):

Let u = <-5, 1>, v = <7, -4>. Find 9u - 6v. <18, 18> <-87, 33> <-36, -18> <-3, -15>

OpenStudy (anonymous):

umm let e check

OpenStudy (anonymous):

yes it is

OpenStudy (kc_kennylau):

9u = <-5\(\times\)9, 1\(\times\)9> = <-45,9> 6v = <7\(\times\)6,-4\(\times\)6> = <42,-24> Understand the above?

OpenStudy (kc_kennylau):

9u-6v = <-45-42,9-(-24)> = <-87,33>

OpenStudy (anonymous):

you just saved me! thank you so much

OpenStudy (kc_kennylau):

The important part is if you get it or not